OpenStudy (zzr0ck3r):

prove that the sequence sin(n*pi/2) diverges using the form (some epsilon)(for all N)(some n)(n>N => |x_n - L| >= epsilon)

OpenStudy (anonymous):

to prove the divergence ! change "=>" by "and"

OpenStudy (zzr0ck3r):

right ty

OpenStudy (anonymous):

what you typed is correct ! just change "=>" by and

OpenStudy (anonymous):

ohh ! I didn't see your reply ! sorry !

OpenStudy (zzr0ck3r):

I know for something like (-1)^n we make a case for n is odd so x_n is -1 and let L >0 and case 2 we let n be even and L<=0. And then show that (x_n-L)>= 1 for some n > all N this is true. I guess I dont understand why we led n odd go with L<0 and n even go with L>= 0.

OpenStudy (anonymous):

Is that a solution ?! because if you're proving the divergence you don't have the right to choose L anymore ! (cause implicitly there is for all L in R before some epsilon !)

OpenStudy (zzr0ck3r):

well it did two cases one for L>=0 and L<0, is that choosing? Im sorry im very new to this stuff and this is my first upper dev class.

OpenStudy (anonymous):

now ! I understand ! just begin with L ! not n ....the cases are for L ! and no problem ! I'm trying to make you understand ! ( well I'm a bad teacher...)

OpenStudy (zzr0ck3r):

yeah I just dont understand why we did not have to show both cases of n odd and both cases for n even.

OpenStudy (anonymous):

you don't have to ! let L>0 so we "have to choose" epsilon=.....as 1 let N be any positive integer !( now we have to choose n) let's take n= 2(N)+1 so obviously n>N |x_n - L|= |-1 - L|=|1 +L| (-1 is a trick to take off | | and...) 1+L>0 then | 1+L|=1+L >=1 the same for L<=0 so n is choosen (using N if we want to !)

OpenStudy (anonymous):

sorry ! Be right back !

OpenStudy (zzr0ck3r):

ahhh I c, we get to choose epsilon and n

OpenStudy (anonymous):

yes ! :) brb

OpenStudy (anonymous):

I'm Back !

OpenStudy (zzr0ck3r):

word. so I can have a case for l > 0 and n = pi/2 + (n-1)(2pi)?

OpenStudy (zzr0ck3r):

and epsilon = 1/2

OpenStudy (anonymous):

for your exercise ?!

OpenStudy (zzr0ck3r):

yeah the sin(npi/2)

OpenStudy (anonymous):

There are many many many ways to do this ! You can do it without using cases !

OpenStudy (zzr0ck3r):

well thats the only way I can think to do it. but I dont know what dto do for the the case when sin(npi/2) = 0

OpenStudy (anonymous):

why are you making cases fo "n" ? you don't have to !

OpenStudy (anonymous):

you can choose epsilon using L....But you have to force it to differ from zero and also satisfies 0< epsilon

OpenStudy (zzr0ck3r):

can you show me?

OpenStudy (anonymous):

Just do it with cases if you can !

OpenStudy (anonymous):

Sorry ! I made a mistake ! With cases is easier !

OpenStudy (anonymous):

Sorry Again ! Just do the same thing As you've done for (-1)^n \[if L >0 \] then let epsilon =1/2 or (1=) let N be any positive integer ! let n= 4(N+1)+3 then sin(n pi/2)=-1 | -1-L|=|1+l=1+L>1/2 foe L<=0.....the same just n = 4(N+1)+1 then |sin( npi/2)-L|=|1-L|=1-L>1/2 (becaus -L>=0) sorry for confusing you @zzr0ck3r

OpenStudy (zzr0ck3r):

nope hah great that is almost exactly what i did just a variation on the n

OpenStudy (anonymous):

Great ! you're learning too fast :) !

OpenStudy (zzr0ck3r):

thank you man, you opened a door for me today and last night. Been waiting for that door to open for about 24 hours:)

OpenStudy (anonymous):

you're welcome ! :) !